# Boter: Bootstrapping Knowledge Selection and Question Answering for Knowledge-based VQA

### v1

@@ Abstract.
Knowledge-based Visual Question Answering (VQA) requires models to incorporate external knowledge to respond to questions about visual content. Previous methods mostly follow the ”retrieve and generate” paradigm. Initially, they utilize a pre-trained retriever to fetch relevant knowledge documents, subsequently employing them to generate answers. While these methods have demonstrated commendable performance in the task, they possess limitations: (1) they employ an independent retriever to acquire knowledge solely based on the similarity between the query and knowledge embeddings, without assessing whether the knowledge document is truly conducive to helping answer the question; (2) they convert the image into text and then conduct retrieval and answering in natural language space, which may not ensure comprehensive acquisition of all image information. To address these limitations, we propose Boter, a novel framework designed to bootstrap knowledge selection and question answering by leveraging the robust multimodal perception capabilities of the Multimodal Large Language Model (MLLM). The framework consists of two modules: Selector and Answerer, where both are initialized by the MLLM and parameter-efficiently finetuned in a simple cycle: find key knowledge in the retrieved knowledge documents using the Selector, and then use them to finetune the Answerer to predict answers; obtain the pseudo-labels of key knowledge documents based on the predictions of the Answerer and weak supervision labels, and then finetune the Selector to select key knowledge; repeat. Our framework significantly enhances the performance of the baseline on the challenging open-domain Knowledge-based VQA benchmark, OK-VQA, achieving a state-of-the-art accuracy of 62.83%. We conduct comprehensive ablations of our framework, including Selector and Answerer ablations, the method of cycle training, varying key knowledge document numbers, the impact of vision information, and so on.
Knowledge-based VQA, Selector and Answerer, Multimodal Large Language Model, Cycle training
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06
Figure 1. Comparison with previous paradigms. (a) Previous KB-VQA paradigms apply captioning tools to transform the image into a visual context. They use a separate pre-trained knowledge retriever to retrieve knowledge and then use them to predict answers in natural language space. However, the top-ranked document contains noise information and misleads the model to answer a wrong answer ”surfing”. (b) The proposed paradigm aims to enhance the utilization of visual information to perform knowledge selection and question-answering consistency. We bootstrap the selector module and answerer module each other in a cycle style: where the Selector chooses key knowledge documents for the Answerer and the Answerer provides pseudo-labels for the Selector. The selector successfully finds key knowledge documents that are low-ranked but informative and the answerer predicts the right answer ”freestyle”.
1. Introduction
Visual Question Answering (VQA) (Antol et al., 2015) is a task that entails answering questions based on an image. Conventional VQA benchmarks (Goyal et al., 2017; Hudson and Manning, 2019) primarily focus on addressing straightforward questions that only necessitate visual perception and recognition. However, in real-world scenarios, many questions require broader world knowledge and common sense for accurate responses. Consequently, the Knowledge-based VQA (KB-VQA) benchmark is introduced to facilitate the development of models suitable for real-world applications. In contrast to previous VQA tasks, KB-VQA presents a more challenging setting where questions grounded in visual content demand answers by referencing external knowledge. Early studies (Wang et al., 2017a, b) offer a pre-defined knowledge base for all questions. More recently, the Outside-Knowledge Visual Question Answering (OK-VQA) (Marino et al., 2019) benchmark is introduced to emphasize the utilization of open-domain world knowledge to answer questions, rather than relying on a predefined knowledge base. This approach simulates a more realistic and human-like interaction with the visual world.
To tackle the knowledge-based Visual Question Answering (VQA) task, a common approach involves retrieving relevant explicit knowledge and leveraging it to answer questions. Existing KB-VQA methods (Marino et al., 2021; Gardères et al., 2020; Lin and Byrne, 2022; Gui et al., 2021; Lin et al., 2022, 2023) often utilize the pre-trained retriever (e.g., DPR (Karpukhin et al., 2020)) to retrieve the related knowledge from diverse explicit knowledge bases, such as ConceptNet (Speer et al., 2017), Wikipedia (Vrandečić and Krötzsch, 2014), and Google Search Corpus (Luo et al., 2021). Subsequently, they utilize the retrieved knowledge to generate answers based on pre-trained language models. In addition to methods utilizing explicit knowledge, some approaches (Gui et al., 2021; Lin et al., 2022; Hu et al., 2022a; Shao et al., 2023) employ GPT-3 (Brown et al., 2020), a large-scale pre-trained language model, to generate implicit textual knowledge relevant to the question. With the explicit and implicit knowledge, the models have achieved good results. In addition to augmenting various knowledge bases, two critical components should be considered to obtain good performance: vision information mining and knowledge selection.
To obtain vision information, many traditional methods transform the image into raw texts, such as captions (Hu et al., 2022a; Lin et al., 2022; Gui et al., 2021), tags (Yang et al., 2022), and OCR (Gao et al., 2022; Lin and Byrne, 2022). They use textual visual information in the process of retrieving and answering, which leads to the underutilization of visual information. On the other hand, to obtain external knowledge, many methods simply use a pre-trained retriever to obtain the relevant knowledge, without considering how to make the answer selector work better with the answer generator. The retrieved documents are only ranked based on the similarity between query and knowledge embeddings, which cannot guarantee that the selected documents can genuinely aid in producing correct answers.
To address the aforementioned limitations, we introduce a novel framework, Boter, designed to select key knowledge documents and answer questions using the Multimodal Large Language Model (MLLM). We aim to enhance the utilization of visual information by leveraging the MLLM’s multimodal perception capabilities.
@@ 1. Introduction
Built upon the MLLM, our framework comprises two modules: a Selector and an Answerer. We train two modules by repeating the following process: the Selector first identifies important knowledge from the candidate knowledge documents retrieved by the pre-trained retriever; then, the Answerer takes the key knowledge documents as the input knowledge and is finetuned to generate the answer; next, we generate pseudo-labels of key knowledge documents according to the Answerer’s predictions and weak supervision labels; finally, we refine the Selector to assess the relevance of retrieved knowledge documents in answering the question. This strategy of cycle training bootstraps the ability of knowledge selection and answer generation consistently, enabling the model to accurately respond to knowledge-intensive questions.
@@
We conduct extensive experiments on the open-domain knowledge-based VQA benchmark (OK-VQA (Marino et al., 2019)) to validate the effectiveness of the proposed framework, where our method largely outperforms the baseline and achieves the state-of-the-art performance of 62.83%, only finetuning 0.16% parameters with LoRA (Hu et al., 2022b). We also conduct comprehensive ablations to validate the impact of different components of the proposed framework, including the Effect of Selector and Answerer, cycle training of the framework, varying the number of key knowledge documents, the impact of vision information, and so on.
@@
  • We introduce a novel framework for consistently selecting key knowledge documents and answering questions using the multimodal language model.
@@
  • We propose a new cycle-training method to bootstrap Selector and Answerer each other, where the Selector chooses key knowledge documents for the Answerer and the Answerer provides pseudo-labels for the Selector.
@@
  • We achieve a state-of-the-art performance of 62.83% on the OK-VQA dataset, surpassing the previous state-of-the-art method. Notably, this improvement is achieved by fine-tuning only 0.16
  • We present a comprehensive analysis to assess the impact of various components within the proposed framework.
2. RELATED WORK
Visual Question Answering (VQA). VQA presents a formidable challenge in visual-language understanding, requiring models to respond to visually-grounded questions. Conventional VQA benchmarks (Goyal et al., 2017; Hudson and Manning, 2019) primarily focus on basic visual perception and reasoning tasks. Numerous studies have achieved promising results on these benchmarks by incorporating: (1) improved visual features (Anderson et al., 2017; Jiang et al., 2020; Shen et al., 2021; Zhang et al., 2021); (2) attention-based cross-modality fusion techniques (Hu et al., 2017; Kim et al., 2018; Li et al., 2019; Yu et al., 2019); (3) the pretrain-finetune paradigm (Tan and Bansal, 2019; Lu et al., 2019; Zhang et al., 2021; Li et al., 2022, 2020; Wang et al., 2022). In addition to these general-purpose VQA tasks, the knowledge-based VQA task (Wang et al., 2017a; Marino et al., 2019) is introduced to advance models’ understanding of the open-ended world. This task leverages world knowledge to answer visual questions, thus enabling models to grasp the broader context of visual queries.
Figure 2. Our Boter framework consists of two modules: a Selector and an Answerer. Selector (left) selects the top-T knowledge documents for the Answerer (right), and the Answerer focuses on important knowledge information to predict answers. Both modules utilize the same frozen visual module to extract image features. We train the fully connected (FC) layer and fine-tune the language model using LoRA, which amounts to only 0.16% of the total parameters. For detailed training procedures of the two modules, refer to Alg. 1. The original knowledge is retrieved using DPR, and for brevity, we omit the retrieval process here (details can be found in Section 3.1).
Knowledge-based VQA. Early benchmarks (Wang et al., 2017a, 2015) provide fixed knowledge bases for the model to retrieve and answer questions. Recently, the open-domain knowledge-based VQA benchmark (OK-VQA (Marino et al., 2019)) is proposed, which provides no extra knowledge bases and the model needs to find proper knowledge bases to help answer questions. Recent studies (Gardères et al., 2020; Wu et al., 2022; Lin and Byrne, 2022; Gui et al., 2021) have explored various open-domain world knowledge sources, such as ConceptNet (Speer et al., 2017), Wikipedia (Vrandečić and Krötzsch, 2014), Google Search Corpus (Luo et al., 2021). They retrieve the relevant knowledge documents from the knowledge bases and integrate them into the answering model to generate predictions. Except for using explicit knowledge, some methods also take GPT-3 (Brown et al., 2020) as an implicit knowledge producer. They either prompt GPT-3 with in-context examples to predict answers directly (Yang et al., 2022; Hu et al., 2022a; Shao et al., 2023), or use GPT-3 to generate answer candidates with evidence serving as textual implicit knowledge bases (Gui et al., 2021; Lin et al., 2022), leading to significant performance improvements.
Different from previous approaches that primarily enhance performance by incorporating more knowledge sources, our focus lies in effectively leveraging visual information and retrieved knowledge. We employ a multimodal large language model to enhance visual understanding and construct a Selector and an Answerer to consistently select key knowledge and answer questions.
@@ 2. RELATED WORK
Multimodal large language model. Recently, multimodal large language models (Li et al., 2023; Alayrac et al., 2022; Liu et al., 2023; Dai et al., 2023) have demonstrated remarkable visual-language understanding and reasoning capabilities, owing to the advancement of larger language models (Brown et al., 2020; Zhang et al., 2022; Touvron et al., 2023; vic, 2023; Touvron et al., 2023). These methods typically consist of a frozen visual encoder (Radford et al., 2021), a visual-language connector (Li et al., 2023), and a large language model (Chung et al., 2022; Zhang et al., 2022; vic, 2023). The models are firstly pre-trained on large-scale visual-text datasets to align visual features to the language embedding space. After pretraining, the large language model can understand the visual details. Then, the model is finetuned to adapt to various visual-language tasks. In this study, we adopt BLIP2, one of the widely used models, as our backbone for understanding visual details and bootstrap knowledge selection and question answering with it.
3. Method
@@
In this section, we first introduce the preliminaries of Knowledge Retrieval and MLLM, which are the foundation of our framework. Then, we present the design of the Selector and Answerer for knowledge selection and question answering on knowledge respectively. Finally, we illustrate the cycle training method of two designed modules.
3.1. Preliminaries
Knowledge Retrieval. Following previous knowledge-based VQA methods (Gao et al., 2022; Lin and Byrne, 2022; Si et al., 2023), we adopt the dense passage retrieval (DPR) (Karpukhin et al., 2020) to retrieve the knowledge documents. We transform the visual image $I_{i}$ into texts through the following steps:
  • The pretrained captioning model Oscar+ (Zhang et al., 2021) is adopted to obtain the image caption $C_{i}$.
  • The pretrained detection model VinVL (Zhang et al., 2021) is adopted to extract the text description of objects and attributes $L_{i}$.
  • The Google OCR (Optical Character Recognition) APIs are adopted to extract the text strings $\{O_{i}\}$ of the image.
Consequently, the visual contexts of the image $I$ can be represented as $V_{i}=(C_{i},L_{i},O_{i})$. We combine the question $Q_{i}$ and the visual contexts $V_{i}$ as a query $q_{i}=(Q_{i},V_{i})$ to retrieve the relevant documents from the knowledge base $D$. DPR encodes the query $q_{i}$ and all the knowledge documents $\mathcal{D}$ into embeddings $\mathbf{q}_{i}$ and $\{\mathbf{d}_{1},\mathbf{d}_{2},...,\mathbf{d}_{n}\}$ with different BERT-like (Devlin et al., 2018) encoders as follows:
(1) $\mathbf{q}_{i}=BERT_{q}(q_{i}),\mathbf{d}_{i}=BERT_{d}(D_{i}),$
@@ 3.1. Preliminaries
We compute the similarity scores $sim(q_{i},D_{j})=\mathbf{q}_{i}^{T}\cdot\mathbf{d}_{j}$ and exploit FAISS (Johnson et al., 2019) to index Top-k related knowledge documents $\mathcal{P}_{i}=\{P_{i,1},P_{i,2},...,P_{i,k}\}$ for $i$-th query.
@@
Multimodal Large Language Model. Recently, multimodal large language models (Li et al., 2023; Alayrac et al., 2022; Liu et al., 2023) have demonstrated significant advancements in visual-language understanding. The architecture of these models typically comprises a frozen image encoder (Dosovitskiy et al., 2020; Fang et al., 2023), a visual-language connector (Li et al., 2023), and a pre-trained language model (Chung et al., 2022; Zhang et al., 2022; vic, 2023). Given an image $I_{i}$, the frozen image encoder outputs a set of visual features $\{\mathbf{h}_{i,1},\mathbf{h}_{i,2},...,\mathbf{h}_{i,m}\}$. Visual-language connector takes extracted visual features as input, and outputs language-aligned visual features $\{\mathbf{v}_{i,1},\mathbf{v}_{i,2},...,\mathbf{v}_{i,l}\}$. These visual features are concatenated with the textual word embeddings, which are fed into the language model for generation. Through pre-training on large-scale image-caption datasets, the visual-language connector can effectively project visual features into the feature space of the Language Large Model (LLM).
In our work, both knowledge selection and question-answering modules adopt BLIP-2 (Li et al., 2023) as the backbone. We freeze the visual encoder and Q-former during training. We train the fully connected layer and use LoRA (Hu et al., 2022b) to finetune the LLM (only finetune 0.16% of total parameters).
3.2. Selector and Answerer
For the knowledge-based VQA task, previous methods (Lin and Byrne, 2022; Gui et al., 2021; Lin et al., 2022) separately use DPR to retrieve relevant knowledge documents from the knowledge base, and then directly use them to help produce the answer for the given question. While these methods have demonstrated the effectiveness of leveraging external knowledge, the separate retriever is not inherently aligned with the answerer. The retrieved knowledge documents are ranked based on similarity scores between the query and knowledge embeddings, which may not always be useful for answering a specific question. Additionally, these methods transform the image into textual information and conduct retrieval and answering in the language space, resulting in suboptimal utilization of visual information.
We adopt a multimodal large language model, BLIP-2 (Li et al., 2023), to enhance the utilization of visual information. We assign BLIP-2 two roles: a Selector and an Answerer. The Selector is responsible for selecting key knowledge from the retrieved documents that truly assist in answering questions, while the Answerer predicts the answer based on the selected knowledge. This integrated approach ensures better alignment between knowledge selection and question answering and enhances the model’s ability to answer questions.

### v3

# Self-Bootstrapped Visual-Language Model for Knowledge Selection and Question Answering

@@ Abstract
While large visual-language models (LVLM) have shown promising results on traditional visual question answering benchmarks, it is still challenging for them to answer complex VQA problems which requires diverse world knowledge. Motivated by the research of retrieval-augmented generation in the field of natural language processing, we use Dense Passage Retrieval (DPR) to retrieve related knowledge to help the model answer questions. However, DPR conduct retrieving in natural language space, which may not ensure comprehensive acquisition of image information. Thus, the retrieved knowledge is not truly conducive to helping answer the question, affecting the performance of the overall system. To address this issue, we propose a novel framework that leverages the visual-language model to select the key knowledge retrieved by DPR and answer questions. The framework consists of two modules: Selector and Answerer, where both are initialized by the LVLM and parameter-efficiently finetuned by self-bootstrapping: find key knowledge in the retrieved knowledge documents using the Selector, and then use them to finetune the Answerer to predict answers; obtain the pseudo-labels of key knowledge documents based on the predictions of the Answerer and weak supervision labels, and then finetune the Selector to select key knowledge; repeat. Our framework significantly enhances the performance of the baseline on the challenging open-domain Knowledge-based VQA benchmark, OK-VQA, achieving a state-of-the-art accuracy of 62.83%. Our code is publicly available at https://github.com/haodongze/Self-KSel-QAns.
Self-Bootstrapped Visual-Language Model for Knowledge Selection and Question Answering
Dongze Hao1,2, Qunbo Wang1, Longteng Guo1, Jie Jiang1, Jing Liu1,2* 1Institute of Automation, Chinese Academy of Sciences, 2School of Artificial Intelligence, University of Chinese Academy of Sciences {haodongze2021,qunbo.wang}@ia.ac.cn, {longteng.guo,jie.jiang,jliu}@nlpr.ia.ac.cn thanks: Jing Liu is the corresponding author
1 Introduction
Recently, there has been an impressive advancement in large visual-language models (LVLM) Li et al. (2023); Alayrac et al. (2022); Liu et al. (2023); Dai et al. (2023). They usually use a mapping network to inject visual features into the semantic space of the large language model Brown et al. (2020); Zhang et al. (2022); Touvron et al. (2023); vic (2023); Touvron et al. (2023) and demonstrate strong capabilities on multimodal perception and reasoning. Thus, they achieve significant progress in conventional visual question answering benchmarks Antol et al. (2015); Goyal et al. (2017); Hudson and Manning (2019) which primarily focus on addressing straightforward questions that only necessitate visual perception and recognition. However, it is still challenging for the LVLMs to answer visual questions which require broader world knowledge and common sense Wang et al. (2017); Marino et al. (2019); Schwenk et al. (2022).
Motivated by the research of retrieval-augmented generation Karpukhin et al. (2020a) in the field of natural language processing, we use Dense Passage Retrieval (DPR) to retrieve related world knowledge to help the model answer questions. However, when using DPR, we need to transform the image into texts to retrieve the related knowledge, which leads to the underutilization of visual information. Thus, the retrieved knowledge may be unfaithful and affects the model performance. To address the issue, we consider the LVLM as the knowledge selector to find helpful knowledge from candidate retrieved knowledge by DPR. Then the selected knowledge is fed into the LVLM to predict the answer.
@@ 1 Introduction
In this paper, we introduce a novel framework where we adopt the large visual-language model to perform knowledge selection and question answering. Our framework comprises two modules: a Selector and an Answerer. We train two modules by repeating the following process: the Selector first identifies important knowledge from the candidate knowledge documents retrieved by the pre-trained retriever; then, the Answerer takes the key knowledge documents as the input knowledge and is finetuned to generate the answer; next, we generate pseudo-labels of key knowledge documents according to the Answerer’s predictions and weak supervision labels; finally, we refine the Selector to assess the relevance of retrieved knowledge documents in answering the question. This strategy of self-bootstrapping enhances the ability of knowledge selection and answer generation consistently, enabling the model to accurately respond to knowledge-intensive questions.
@@
We conduct extensive experiments on the open-domain knowledge-based VQA benchmark (OK-VQA Marino et al. (2019)) to validate the effectiveness of the proposed framework, where our method largely outperforms the baseline and achieves the state-of-the-art performance of 62.83%, only finetuning 0.16% parameters with LoRA Hu et al. (2022a). We also conduct comprehensive ablations to validate the impact of different components of the proposed framework, including the Effect of Selector and Answerer, cycle training of the framework, varying the number of key knowledge documents, and so on.
@@
  • We introduce a novel framework that leverages the large visual-language model to select key knowledge and use them to answer questions, respectively.
@@
  • We propose a new self-bootstrap learning method to train the Selector and Answerer, where the Selector chooses key knowledge documents for the Answerer and the Answerer provides pseudo-labels for the Selector.
@@
  • We achieve a state-of-the-art performance of 62.83% on the OK-VQA dataset, surpassing the previous state-of-the-art method. Notably, this improvement is achieved by fine-tuning only 0.16% of parameters using LoRA.
2 Related work
Large Visual-Language Models.
@@ 2 Related work
Recently, large visual-language models Li et al. (2023); Alayrac et al. (2022); Liu et al. (2023); Dai et al. (2023) have demonstrated remarkable visual-language understanding and reasoning capabilities, owing to the advancement of larger language models Brown et al. (2020); Zhang et al. (2022); Touvron et al. (2023); vic (2023); Touvron et al. (2023). These methods typically consist of a frozen visual encoder Radford et al. (2021), a visual-language connector Li et al. (2023), and a large language model Chung et al. (2022); Zhang et al. (2022); vic (2023). The models are firstly pre-trained on large-scale visual-text datasets to align visual features to the language embedding space. After pretraining, the large language model can understand the visual details. Then, the model is finetuned to adapt to various visual-language tasks. In this study, we adopt BLIP2, one of the widely used models, as our backbone for bootstrapping knowledge selection and question answering with it.
Figure 1: Our framework consists of two modules: a Selector and an Answerer. Selector (left) selects the top-T knowledge documents for the Answerer (right), and the Answerer focuses on important knowledge information to predict answers. Both modules utilize the same frozen visual module to extract image features. We train the fully connected (FC) layer and fine-tune the language model using LoRA, which amounts to only 0.16% of the total parameters. For detailed training procedures of the two modules, refer to Alg. 1. The original knowledge is retrieved using DPR, and for brevity, we omit the retrieval process here (details can be found in Section 3.1).
Knowledge-based VQA.
Conventional VQA benchmarks Goyal et al. (2017); Hudson and Manning (2019) primarily focus on basic visual perception and reasoning tasks and numerous studies have achieved promising results on these benchmarks Anderson et al. (2017); Zhang et al. (2021); Tan and Bansal (2019); Lu et al. (2019); Li et al. (2022); Wang et al. (2022). Different from them, the knowledge-based VQA task Wang et al. (2017); Marino et al. (2019); Schwenk et al. (2022) requires models to incorporate diverse world knowledge to respond to questions about visual content, which is more challenging. Recent studies Gardères et al. (2020); Wu et al. (2022); Lin and Byrne (2022); Gui et al. (2021) have explored various open-domain world knowledge sources, such as ConceptNet Speer et al. (2017), Wikipedia Vrandečić and Krötzsch (2014), Google Search Corpus Luo et al. (2021). They retrieve the relevant knowledge documents from the knowledge bases and integrate them into the answering model to generate predictions. Except for using explicit knowledge, some methods also take GPT-3 Brown et al. (2020) as an implicit knowledge producer. They either prompt GPT-3 with in-context examples to predict answers directly Yang et al. (2022); Hu et al. (2022b); Shao et al. (2023), or use GPT-3 to generate answer candidates with evidence serving as textual implicit knowledge bases Gui et al. (2021); Lin et al. (2022), leading to significant performance improvements. Different from these approaches, we employ a large visual-language model to select key retrieved knowledge and reason on the knowledge to answer questions.
3 Method
@@
In this section, we first introduce the preliminaries of Knowledge Retrieval and LVLM, which are the foundation of our framework. Then, we present the design of the Selector and Answerer for knowledge selection and question answering on knowledge respectively. Finally, we illustrate the self-bootstrap training method of two designed modules.
3.1 Preliminaries
Knowledge Retrieval.
@@ 3.1 Preliminaries
We adopt the Dense Passage Retrieval (DPR) Karpukhin et al. (2020b) to retrieve the knowledge documents. We transform the image into raw texts composed of captions, objects, attributes, and OCR (Optical Character Recognition). Then we compute the similarity scores between the query and knowledge documents $sim(q_{i},D_{j})=\mathbf{q}_{i}^{T}\cdot\mathbf{d}_{j}$ and exploit FAISS Johnson et al. (2019) to index Top-k related knowledge documents $\mathcal{P}_{i}=\{P_{i,1},P_{i,2},...,P_{i,k}\}$ for $i$-th query.
Large Visual-Language Model.
@@
In our work, both knowledge selection and question-answering modules adopt BLIP-2 Li et al. (2023) as the backbone. The architecture of BLIP-2 comprises a frozen image encoder Dosovitskiy et al. (2020); Fang et al. (2023), a Q-Former Li et al. (2023), and a pre-trained language model Chung et al. (2022). Given an image $I_{i}$, the frozen image encoder outputs a set of visual features $\{\mathbf{h}_{i,1},\mathbf{h}_{i,2},...,\mathbf{h}_{i,m}\}$. Q-Former takes extracted visual features as input, and outputs language-aligned visual features $\{\mathbf{v}_{i,1},\mathbf{v}_{i,2},...,\mathbf{v}_{i,l}\}$. These visual features are concatenated with the textual word embeddings, which are fed into the language model for generation. Through pre-training on large-scale image-caption datasets, Q-Former can effectively project visual features into the feature space of the Language Large Model (LLM). We freeze the visual encoder and Q-former during training. We train the fully connected layer and use LoRA Hu et al. (2022a) to finetune the LLM (only finetune 0.16% of total parameters).
3.2 Selector and Answerer
Selector.